OpenStudy (anonymous):

5(2x - 7) = 33 - 3(6x + 4) type out in words the steps you took to solve the equation

OpenStudy (anonymous):

multiply the brackets with 5 and 3 respectively 10x-35=33-18x-12 simplify 28x=56 solve for x divide by 28 x=2

OpenStudy (anonymous):

did u get it ??

OpenStudy (anonymous):

yeah thank you!!

OpenStudy (anonymous):

hashi has right idea :) i'll dbl check

OpenStudy (anonymous):

my pleasure

OpenStudy (anonymous):

@timo86m: its hashir

OpenStudy (anonymous):

x=2 yes :)

OpenStudy (anonymous):

it says it should have the minimum of 7 steps

OpenStudy (anonymous):

lol sorry my keyboard is no good. Oh wait OS is not good.

OpenStudy (anonymous):